OpenStudy (anonymous):

Given that f(x) = x2 + 3x + 6 and g(x) = the quantity of x plus two, over three, solve for f(g(x)) when x = 1.

OpenStudy (anonymous):

Let's first put g(x) into its algebraic representation. "The quantity of.." part immediately tells you that you're going to have parenthesis, and that (x+2) is one entity. Then, you need to divide it by 3. So g(x) = (x+2)/3. If the question didn't include "the quantity of..", then g(x) would be ambiguous. We could interpret it to be x + 2/3, which is not correct. Anyway, f(g(x)) says to take every 'x' in f(x), and replace them with the entire function g(x). So, \[ f(x) = \left( \frac{x+2}{3} \right)^2 + 3\left(\frac{x+2}{3}\right) + 6. \] You can simplify it you want.

OpenStudy (anonymous):

So i plug in 1?

OpenStudy (anonymous):

If i plug in 1 then for g(x) i got 1 and for f(x) i got 10 :o

OpenStudy (anonymous):

Oh right it wants it 1. So you don't even need to simplify! That would just waste time. So yeah f(g(1)) = (3/3)^2 + 3(1+2)/3 + 6 = 1 + 3 + 6 = 10. That's correct.

OpenStudy (anonymous):

but the choices are A.) -1 B.) 0 C.) 3 D.) 4 So what would i do next?